OpenStudy (queenb101):

.list three different ways to write 5^11 as the product of two prowers .Explain why all three of your expressions are equal to 5^11

OpenStudy (queenb101):

plz help

Vocaloid (vocaloid):

when you multiply powers, keep the base and add the exponent so basically, 5^___ * 5^_____ = 5^11 where two numbers in the blanks add up to 11

Vocaloid (vocaloid):

ex: 5 + 6 = 11 so 5^5 * 5^6 = 11 now, that's one way. find two other ways.

OpenStudy (queenb101):

ok

OpenStudy (queenb101):

so um let see another way is 0 x 11 =11 5^0 x 5^11